# LLM-Mixer: Multiscale Mixing in LLMs for Time Series Forecasting

### v1

@@ Abstract
Time series forecasting remains a challenging task, particularly in the context of complex multiscale temporal patterns. This study presents LLM-Mixer, a framework that improves forecasting accuracy through the combination of multiscale time-series decomposition with pre-trained LLMs (Large Language Models). LLM-Mixer captures both short-term fluctuations and long-term trends by decomposing the data into multiple temporal resolutions and processing them with a frozen LLM, guided by a textual prompt specifically designed for time-series data. Extensive experiments conducted on multivariate and univariate datasets demonstrate that LLM-Mixer achieves competitive performance, outperforming recent state-of-the-art models across various forecasting horizons. This work highlights the potential of combining multiscale analysis and LLMs for effective and scalable time-series forecasting. Code is available at: https://github.com/Kowsher/LLMMixer.
@@
Md Kowsher1, Md. Shohanur Islam Sobuj2, Nusrat Jahan Prottasha1, E. Alejandro Alanis3, Ozlem Ozmen Garibay1, Niloofar Yousefi1 1University of Central Florida, USA 2Anymate Me, Germany, 3Microsoft, USA
@@ 1 Introduction & Related Work
Time series forecasting is essential in numerous fields, including finance Zhang et al. (2024), energy management Martín et al. (2010), healthcare Morid et al. (2023), climate science Mudelsee (2019), and industrial operations Wang et al. (2020). Traditional forecasting models, such as AutoRegressive Integrated Moving Average (ARIMA) (Box et al., 2015) and exponential smoothing techniques (Hyndman, 2018), are widely used for straightforward predictive tasks. However, these models assume stationarity and linearity, which limit their effectiveness when applied to complex, nonlinear, and multivariate time series often found in real-world scenarios Cheng et al. (2015). The advent of deep learning has significantly advanced time series forecasting. CNNs Wang et al. (2023); Tang et al. (2020); Kirisci and Cagcag Yolcu (2022) have been utilized for capturing temporal patterns, while RNNs Siami-Namini et al. (2019); Zhang et al. (2019); Karim et al. (2019) are adept at modeling temporal state transitions. However, both CNNs and RNNs have limitations in capturing long-term dependencies Wang et al. (2024); Tang et al. (2021); Zhu et al. (2023). Recently, Transformer architectures Vaswani et al. (2017) have demonstrated strong capabilities in handling both local and long-range dependencies, making them suitable for time series forecasting (Liu et al., 2024b; Nie et al., 2022; Woo et al., 2022).
@@
In parallel, pre-trained LLMs such as GPT-3 (Brown, 2020), GPT-4 (Achiam et al., 2023), and LLaMA (Touvron et al., 2023) have achieved remarkable generalization in natural language processing tasks Friha et al. (2024) due to capabilities of few-shot or zero-shot transfer learning Brown (2020), multimodal knowledge Jia et al. (2024) and reasoning Liu et al. (2024a). These models are now being applied across various fields, including computer vision Bendou et al. (2024), healthcare Gebreab et al. (2024), and finance Zhao et al. (2024). Recently, a few studies have explored using LLMs for time series forecasting due to their impressive capabilities Jin et al. (2024, 2023); Gruver et al. (2023). However, adapting LLMs to time series data presents challenges because there are significant differences between token-based text data and continuous time series data Morales-García et al. (2024). LLMs are built to handle discrete tokens, which limits their ability to capture the continuous and often irregular patterns found in time series data. Additionally, time series data has multiple time scales, from short-term fluctuations to long-term trends, making it difficult for traditional LLMs to capture all these patterns at once. LLMs typically process fixed-length sequences, which means they may only capture short-term dependencies if the sequence length (i.e., the window of time steps) is small. However, extending the sequence length to capture long-term trends increases computational costs and may dilute the model’s ability to focus on short-term fluctuations within the same sequence. Previous studies using LLMs on time series data have mostly fed the original sequence directly into a frozen LLM, making it hard for the model to fully understand these sequences.
@@ 2 LLM Mixer
Preliminaries: In multivariate time series forecasting, we are given historical data $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}\in\mathbb{R}^{T\times M}$, where $T$ is the number of time steps and $M$ is the number of variables. The goal is to predict the future values for the next $K$ time steps, denoted as $\mathbf{Y}=\{\mathbf{x}_{T+1},\dots,\mathbf{x}_{T+K}\}\in\mathbb{R}^{K\times M}$. For convenience, let $\mathbf{X}_{t,:}$ represent the data at time step $t$, and $\mathbf{X}_{:,m}$ represent the full time series for variable $m\in M$.
@@
We then use stacked Past-Decomposable-Mixing (PDM) blocks from Jiménez-Navarro et al. (2023) to mix past information across different scales. PDM works by breaking down complex time series data into separate seasonal and trend components at multiple scales, allowing for targeted processing of each component by following Wu et al. (2021). For the $l$-th layer, PDM is defined as
@@
Methods
LLM-Mixer
LLM-Mixer
ETTh1 0.420 0.433 0.434 0.454 0.421 0.437 0.447 0.440 0.454 0.447 0.446 0.434 0.461 0.457 0.516 0.484 0.495 0.450 0.541 0.507 0.458 0.450 0.529 0.522
ETTh2 0.345 0.389 0.349 0.395 0.349 0.391 0.364 0.395 0.383 0.407 0.374 0.398 0.563 0.519 0.391 0.411 0.414 0.427 0.611 0.550 0.414 0.427 0.942 0.684
ETTm1 0.367 0.387 0.374 0.396 0.367 0.388 0.381 0.395 0.407 0.410 0.414 0.407 0.404 0.408 0.406 0.407 0.400 0.406 0.419 0.419 0.400 0.406 0.513 0.495
ETTm2 0.265 0.320 0.272 0.325 0.264 0.319 0.275 0.323 0.288 0.332 0.286 0.327 0.354 0.402 0.290 0.334 0.291 0.333 0.358 0.404 0.291 0.333 0.757 0.610
Weather 0.235 0.264 0.252 0.269 0.235 0.264 0.240 0.271 0.258 0.278 0.272 0.291 0.265 0.315 0.265 0.285 0.251 0.294 0.271 0.320 0.259 0.287 0.264 0.320
Electricity 0.171 0.253 0.174 0.273 0.173 0.261 0.182 0.272 0.178 0.270 0.219 0.298 0.225 0.319 0.216 0.318 0.193 0.304 0.251 0.344 0.192 0.295 0.244 0.334
Traffic 0.414 0.265 0.433 0.282 0.415 0.279 0.484 0.297 0.428 0.282 0.626 0.378 0.625 0.383 0.529 0.341 0.620 0.336 0.760 0.473 0.620 0.336 0.667 0.426
Table 1: Long-term multivariate forecasting results in terms of MSE and MAE, the lower the better. See Table 5 in the Appendix for the full results. Red: the best, Blue: the second best.
Methods
LLM-Mixer
LLM-Mixer
PEMS03 0.112 0.220 0.116 0.226 0.120 0.230 0.116 0.225 0.113 0.221 0.495 0.472 0.180 0.291 0.169 0.281 0.326 0.419 0.147 0.248 0.278 0.375 0.114 0.224
PEMS04 0.097 0.205 0.104 0.211 0.100 0.209 0.098 0.207 0.111 0.221 0.526 0.491 0.195 0.307 0.209 0.314 0.353 0.475 0.129 0.245 0.329 0.395 0.119 0.234
PEMS07 0.102 0.210 0.107 0.217 0.104 0.212 0.104 0.190 0.101 0.204 0.504 0.478 0.213 0.303 0.215 0.326 0.355 0.499 0.129 0.245 0.529 0.387 0.119 0.234
PEMS08 0.130 0.220 0.132 0.224 0.131 0.223 0.132 0.224 0.119 0.226 0.503 0.501 0.243 0.353 0.225 0.357 0.360 0.470 0.206 0.329 0.368 0.393 0.118 0.212
Table 2: Short-term multivariate forecasting results. See Table 6 in Appendix for the full results
Methods
LLM-Mixer
LLM-Mixer
ETTh1 0.068 0.206 0.071 0.207 0.133 0.287 0.071 0.208 0.104 0.247 0.111 0.257 0.108 0.255 0.106 0.252 0.199 0.377 0.345 0.513
ETTh2 0.177 0.332 0.178 0.334 0.204 0.356 0.179 0.335 0.198 0.350 0.205 0.350 0.238 0.384 0.218 0.364 0.243 0.400 0.252 0.408
ETTm1 0.045 0.161 0.047 0.163 0.053 0.167 0.048 0.163 0.054 0.168 0.069 0.202 0.070 0.203 0.081 0.221 0.281 0.441 0.231 0.381
ETTm2 0.109 0.243 0.110 0.243 0.114 0.250 0.110 0.246 0.112 0.248 0.119 0.262 0.135 0.279 0.130 0.271 0.150 0.295 0.130 0.277
Table 3: Long-term univariate forecasting results. See Table 7 in Appendix for the full results.
3 Experiments
We evaluate our LLM-Mixer on several datasets commonly used for benchmarking long-term and short-term multivariate forecasting and compared with SOTA baselines. For long-term forecasting, we use the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) from Zhou et al. (2021), as well as Weather, Electricity, and Traffic datasets from Zeng et al. (2023). For short-term forecasting, we use the PeMS dataset Chen et al. (2001), which consists of four public traffic network datasets (PEMS03, PEMS04, PEMS07, PEMS08) with time series collected at various frequencies. We used RoBERTa-base Liu et al. (2019) as a medium-sized language model and LLaMA2-7B Touvron et al. (2023) as a large language model for the backbone of our framework. Details about the dataset, baselines, hyperparameters, and implementation specifics are provided in Appendices B, C, E and D, respectively.
Multivariate forecasting results: LLM-Mixer demonstrates competitive performance in multivariate long forecasting, as shown in Table 1. Averaged over four forecasting horizons (96, 192, 384, and 720), LLM-Mixer achieves consistently low MSE and MAE values across most datasets, particularly excelling on ETTh1, ETTh2, and Electricity. Compared to other models like TIME-LLM, TimeMixer, and PatchTST, LLM-Mixer performs favorably, showing that its design effectively captures both short and long-term dependencies. Notably, LLM-Mixer also exhibits robustness on challenging datasets such as Traffic, where it outperforms several baseline models. These results highlight the efficacy of LLM-Mixer in handling complex temporal patterns over extended horizons.
Short-term forecasting results: In Table 2, we present the short-term multivariate forecasting results, averaged across four forecasting horizons: 12, 24, 48, and 96 time steps. Our proposed model consistently achieves low MSE and MAE values across the PEMS datasets, indicating strong short-term predictive performance. Specifically, LLM-Mixer demonstrates competitive accuracy on PEMS03, PEMS04, and PEMS07, outperforming several baseline models, including TIME-LLM, TimeMixer, and PatchTST. Additionally, LLM-Mixer shows robustness on PEMS08, where it delivers superior results compared to iTransformer and DLinear. These results emphasize the effectiveness of LLM-Mixer in capturing essential temporal dynamics for short-horizon forecasting tasks.
Univariate forecasting results: Table 3 presents the univariate long forecasting results on the ETT benchmark, averaged over horizons of 96, 192, 384, and 720 time steps. Our proposed method, LLM-Mixer, achieves the lowest MSE and MAE values across all datasets, consistently outperforming other methods like Linear, NLinear, and FEDformer. LLM-Mixer demonstrates superior accuracy, particularly on most of the datasets. These results confirm LLM-Mixer’s effectiveness in capturing complex temporal dependencies, solidifying its capability for univariate long-term forecasting.
3.1 Ablation Study
Effect of Downsampling on NTK Distance: To assess the impact of downsampling levels on model learning dynamics, we conducted an ablation study using the Neural Tangent Kernel (NTK) Jacot et al. (2018). We used a reference model with 10 downsampling levels and computed the NTK for 300 sample pairs from the each training and test sets. The NTK was then compared across models with fewer downsampling levels by calculating the Frobenius norm of the difference between each model’s NTK and that of the reference model. Let $\mathbf{K}_{10}$ denote the NTK of the reference model with 10 downsampling levels. For each reduced downsampling level $\tau\in\{1,2,\dots,9\}$, we computed the NTK, denoted $\mathbf{K}_{\tau}$, and calculated the NTK distance by $d_{\text{NTK}}(\tau)=\|\mathbf{K}_{10}-\mathbf{K}_{\tau}\|_{F}$, where $\|\cdot\|_{F}$ represents the Frobenius norm. Our results, shown in Figure 2 (Left), indicate that as the number of downsampling levels decreases, the NTK distance increases, with the highest distance observed when only one downsampling level is used. This suggests that reducing downsampling progressively alters the model’s learning capacity, indicating that multiscale information plays a crucial role in preserving the original NTK structure. See Appendix A for details on NTK and related experiments.
Multi-scale Mixing by Pooling: We conducted an ablation study to explore the effects of various Multi-scale Mixing techniques. The techniques examined were Min, Max, Avg, and L2, each applying a unique method for aggregating downsampling information across scales. Figure 2 (Right) presents the MSE for each downsampling method across the different datasets. Notably, the average pooling consistently yielded lower MSE, suggesting that this method is better suited for capturing multi-scale dependencies in the data.
4 Conclusion
This work introduces LLM-Mixer, a novel framework that combines multiscale time series decomposition with pre-trained LLMs for improved forecasting. By leveraging multiple temporal resolutions, LLM-Mixer effectively captures both short-term and long-term patterns, enhancing the model’s predictive accuracy. Our experiments demonstrate that LLM-Mixer achieves competitive performance across various datasets, outperforming recent state-of-the-art methods.
5 Limitations
While LLM-Mixer improves forecasting accuracy, it has some limitations. Using pre-trained language models can be computationally demanding, which may limit their application in real-time or large-scale settings. The success of the model also depends on the quality of the prompts used, which could be further optimized. Additionally, testing the model on more diverse datasets would help confirm its reliability across different fields.
6 Ethical Considerations
In this study, we created an LLM-Mixer to improve time-series forecasting using publicly available datasets and pre-trained language models. Our research is purely for academic purposes, and we do not foresee any direct ethical issues. However, because the model can be applied in areas such as finance or healthcare, it is important to ensure that it is used responsibly.
Appendix A NTK
Definition 1 (Neural Tangent Kernel - NTK): Let $\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})$ denote the Neural Tangent Kernel (NTK) of a model. The NTK is defined as the inner product of the gradients of the model outputs with respect to the parameters $\boldsymbol{\theta}$. Specifically, for inputs $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}$, the kernel is given by:
$\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})=\nabla_{\boldsymbol{\theta}}\phi_{P}(\mathbf{x};\boldsymbol{\theta})^{\top}\nabla_{\boldsymbol{\theta}}\phi_{P}(\mathbf{x}^{\prime};\boldsymbol{\theta}),$
where $\nabla_{\boldsymbol{\theta}}\phi_{P}(\mathbf{x};\boldsymbol{\theta})$ denotes the gradient of the model output $\phi_{P}(\mathbf{x};\boldsymbol{\theta})$ with respect to the parameters $\boldsymbol{\theta}$.
Let $\mathbf{K}_{10}\in\mathbb{R}^{n\times n}$ represent the NTK matrix for a reference model with 10 downsampling levels, where $n$ is the number of data samples considered. For each reduced downsampling level $\tau\in\{1,2,\dots,9\}$, we define the corresponding NTK as $\mathbf{K}_{\tau}\in\mathbb{R}^{n\times n}$, computed from the gradient inner product structure of the model at that specific downsampling level.
To quantify the divergence of each reduced NTK $\mathbf{K}_{\tau}$ from the reference NTK $\mathbf{K}_{10}$, we define the NTK distance $d_{\text{NTK}}(\tau)$ as the Frobenius norm of their difference:
$d_{\text{NTK}}(\tau)=\|\mathbf{K}_{10}-\mathbf{K}_{\tau}\|_{F}.$
The Frobenius norm $\|\mathbf{A}\|_{F}$ for any matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ is defined as:
$\|\mathbf{A}\|_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}|A_{ij}|^{2}},$
where $A_{ij}$ denotes the $(i,j)$-th entry of the matrix $\mathbf{A}$.
Thus, the NTK distance $d_{\text{NTK}}(\tau)$ between $\mathbf{K}_{10}$ and $\mathbf{K}_{\tau}$ is given by:
$d_{\text{NTK}}(\tau)=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}|(\mathbf{K}_{10})_{ij}-(\mathbf{K}_{\tau})_{ij}|^{2}}.$
This distance measure captures the total element-wise discrepancy between the NTK matrices of the reference and reduced models, where a higher $d_{\text{NTK}}(\tau)$ indicates a greater deviation in kernel behavior as the downsampling level $\tau$ decreases from 10. The NTK distances for various downsampling levels are as follows:
  • $\tau=10$: $d_{\text{NTK}}(10)=0.0$
  • $\tau=9$: $d_{\text{NTK}}(9)=0.127$
  • $\tau=8$: $d_{\text{NTK}}(8)=0.144$
  • $\tau=7$: $d_{\text{NTK}}(7)=0.215$
  • $\tau=6$: $d_{\text{NTK}}(6)=0.354$
  • $\tau=5$: $d_{\text{NTK}}(5)=0.659$
  • $\tau=4$: $d_{\text{NTK}}(4)=0.708$
  • $\tau=3$: $d_{\text{NTK}}(3)=1.096$
  • $\tau=2$: $d_{\text{NTK}}(2)=1.073$
  • $\tau=1$: $d_{\text{NTK}}(1)=1.167$
This increasing trend in NTK distance indicates a significant change in learning dynamics when fewer downsampling levels are used, suggesting the importance of preserving multiscale structures to maintain effective model performance.
In Figure 3, we visualize the NTK of the reference model across different downsampling levels $\tau$ and the normalized absolute differences.
Appendix B Datasets Description
Tasks Dataset Dim Series Length Dataset Size Frequency Forecastability* Information
Long-term Forecasting ETTm1 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min 0.46 Temperature
ETTm2 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min 0.55 Temperature
ETTh1 7 {96, 192, 336, 720} (8545, 2881, 2881) 15 min 0.38 Temperature
ETTh2 7 {96, 192, 336, 720} (8545, 2881, 2881) 15 min 0.45 Temperature
Electricity 321 {96, 192, 336, 720} (18317, 2633, 5261) Hourly 0.77 Electricity
Traffic 862 {96, 192, 336, 720} (12185, 1757, 3509) Hourly 0.68 Transportation
Weather 21 {96, 192, 336, 720} (36792, 5271, 10540) 10 min 0.75 Weather
Short-term Forecasting PEMS03 358 12 (15617, 5135, 5135) 5min 0.65 Transportation
PEMS04 307 12 (10172, 3375, 3375) 5min 0.45 Transportation
PEMS07 883 12 (16911, 5622, 5622) 5min 0.58 Transportation
PEMS08 170 12 (10690, 3548, 265) 5min 0.52 Transportation
Table 4: Dataset detailed descriptions. The dataset size is organized in (Train, Validation, Test).
The used dataset has been described in Table 4
Appendix C Baselines
We compare our model with well-established time-series forecasting baselines such as TimeMixer Wang et al. (2024), iTransformer Liu et al. (2024b), TimeLLM Jin et al. (2024), RLinear Li et al. (2024), SCINet LIU et al. (2022), TimesNet Wu et al. (2022), TiDE Das et al. (2023), DLinear Zeng et al. (2023), PatchTST Nie et al. (2022), FEDformer Zhou et al. (2022), Stationary Liu et al. (2022), ESTformer Woo et al. (2022), LightTS Campos et al. (2023), and Autoformer Chen et al. (2021). Additionally, we include LLM-based systems such as TimeLLM Jin et al. (2024) and GPT2TS Zhou et al. (2023). For multivariate time series forecasting, we follow the setup of Wang et al. (2024). For short-term forecasting, we adopt the settings from Liu et al. (2024b), and for univariate forecasting, we adhere to the approach in Zeng et al. (2023).
Appendix D Implementation Details
All experiments in this work are implemented using PyTorch. We utilize the Hugging Face library for the LLM model. Experiments were conducted on an NVIDIA H100 GPU with 80 GB of RAM.
Appendix E Hyperparameters
In this work, all long-term experiments use a look-back window of 96 for the RoBERTa-base model, and 192 for the LLaMA2-7B model. For short-term experiments, we use windows of 24 and 48. The RoBERTa-based experiments are conducted for 50 epochs, while LLaMA2 experiments are run for 10 epochs. We employ a batch size of 64 for RoBERTa, and a batch size of 8 with gradient accumulation of 4 for LLaMA2. All experiments utilize the ADAM optimizer with default hyperparameter settings for $(\beta_{1},\beta_{2})=(0.9,0.999)$, and a learning rate of 0.0001.
Appendix F Full Results
The full results are provided in Table 5, 6 and 7.

### v2

@@ Abstract
Time series forecasting is a challenging task, especially when dealing with data that contains both short-term variations and long-term trends. In this study, we introduce LLM-Mixer, a novel framework that combines multiscale time-series decomposition with the power of pre-trained Large Language Models (LLMs). LLM-Mixer breaks down time-series data into multiple temporal resolutions using downsampling and processes these multiscale representations with a frozen LLM, guided by a carefully designed text prompt that encodes information about the dataset’s features and structure. To understand the role of downsampling, we conduct a detailed analysis using Neural Tangent Kernel (NTK) distance, showing that incorporating multiple scales improves the model’s learning dynamics. We evaluate LLM-Mixer across a diverse set of forecasting tasks, including long-term multivariate, short-term multivariate, and long-term univariate scenarios. Experimental results demonstrate that LLM-Mixer achieves competitive performance compared to recent state-of-the-art models across various forecasting horizons. Code is available at: https://github.com/Kowsher/LLMMixer
@@
[…]
@@ 1 Introduction & Related Work
Time series forecasting is essential in numerous fields, including finance Zhang et al. (2024), energy management Martín et al. (2010), healthcare Morid et al. (2023), climate science Mudelsee (2019), and industrial operations Wang et al. (2020). Traditional forecasting models, such as AutoRegressive Integrated Moving Average (ARIMA) Box et al. (2015) and exponential smoothing techniques Hyndman (2018), are widely used for straightforward predictive tasks. However, these models assume stationarity and linearity, which limit their effectiveness when applied to complex, nonlinear, and multivariate time series often found in real-world scenarios Cheng et al. (2015). The advent of deep learning has significantly advanced time series forecasting. CNNs Wang et al. (2023); Tang et al. (2020); Kirisci and Cagcag Yolcu (2022) have been utilized for capturing temporal patterns, while RNNs Siami-Namini et al. (2019); Zhang et al. (2019); Karim et al. (2019) are adept at modeling temporal state transitions. However, both CNNs and RNNs have limitations in capturing long-term dependencies Wang et al. (2024); Tang et al. (2021); Zhu et al. (2023). Recently, Transformer architectures Vaswani et al. (2017) have demonstrated strong capabilities in handling both local and long-range dependencies, making them suitable for time series forecasting Liu et al. (2024b); Nie et al. (2022); Woo et al. (2022).
@@
In parallel, pre-trained LLMs such as GPT-3 Brown (2020), GPT-4 Achiam et al. (2023), and LLaMA Touvron et al. (2023) have achieved remarkable generalization in natural language processing tasks Friha et al. (2024) due to capabilities of few-shot or zero-shot transfer learning Brown (2020), multimodal knowledge Jia et al. (2024) and reasoning Liu et al. (2024a). These models are now being applied across various fields, including computer vision Bendou et al. (2024), healthcare Gebreab et al. (2024), and finance Zhao et al. (2024). Recently, a few studies have explored using LLMs for time series forecasting due to their impressive capabilities Jin et al. (2024, 2023); Gruver et al. (2023). However, adapting LLMs to time series data presents challenges because there are significant differences between token-based text data and continuous time series data Morales-García et al. (2024). LLMs are built to handle discrete tokens, which limits their ability to capture the continuous and often irregular patterns found in time series data. Additionally, time series data has multiple time scales, from short-term fluctuations to long-term trends, making it difficult for traditional LLMs to capture all these patterns at once. LLMs typically process fixed-length sequences, which means they may only capture short-term dependencies if the sequence length (i.e., the window of time steps) is small. However, extending the sequence length to capture long-term trends increases computational costs and may dilute the model’s ability to focus on short-term fluctuations within the same sequence. Previous studies using LLMs on time series data have mostly fed the original or a single sequence directly into a frozen LLM, making it hard for the model to fully understand these sequences Jin et al. (2024, 2023); Gruver et al. (2023).
@@ 2 LLM Mixer
Preliminaries: In multivariate time series forecasting, we are given historical data $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{T}\}\in\mathbb{R}^{T\times M}$, where $T$ is the number of time steps and $M$ is the number of features. The goal is to predict the future values for the next $K$ time steps, denoted as $\mathbf{Y}=\{\mathbf{x}_{T+1},\dots,\mathbf{x}_{T+K}\}\in\mathbb{R}^{K\times M}$. For convenience, let $\mathbf{X}_{t,:}$ represent the data at time step $t$, and $\mathbf{X}_{:,m}$ represent the full time series for variable $m\in M$.
@@
We then use stacked Past-Decomposable-Mixing (PDM) blocks by following the framework from Wang et al. (2024); Jiménez-Navarro et al. (2023) to mix past information across different scales. PDM works by breaking down complex time series data into separate seasonal and trend components at multiple scales, allowing for targeted processing of each component by using the framework from Wang et al. (2024); Wu et al. (2021). For the $l$-th layer, PDM is defined as